# Hamiltonian stochastic Lie systems and applications

### v2

@@ 2 Stochastic differential equations and stochastic Lie systems
where the integrals appearing above are Stratonovich integrals. It is relevant to understand that stochastic differential equations can still be understood in the so-called Itô way, in which the general solution is also of the form (2.4), but the integrals are assumed to be Itô integrals, which is different. If one approach or the other is used, depends on the applications to be developed.
Unless otherwise stated, every stochastic differential equation (2.3) is assumed to be in the Stratonovich sense. This is motivated by the Malliavin transfer principle [42, 45], which suggests that the obtained theory will retain the standard differential theory, although it may not be always this way. In other words, Malliavin transfer principle states that “A formula which is true in the deterministic context and which has a meaning via Stratonovitch stochastic calculus, is still valid, but only almost surely” [42].
Despite this general idea, stochastic differential equations appear in the Itô sense in many applications [28, 48, 55],. Hence, a manner to deal with such stochastic differential equations will be studied in this work. Moreover, the analysis of such stochastic differential equations show some differences appearing in the stochastic formalism with respect to the deterministic theory of Lie systems, which enriches the theory.
Although every Stratonovich stochastic differential equation is equivalent to another Itô stochastic differential equation, the form of both is different. Hence, one has to take care of the method employed to study a stochastic differential equation [29]. More exactly, assume that (2.2) has coefficients that do not depend on ${B}^{2},\ldots,{B}^{\ell}$, then the Itô differential equation admitting the same solutions reads (see [11, p. 137] for details)
@@
$\delta I=\left(-\sigma^{2}(t)I^{3}+\left(-\beta+\frac{3N\sigma^{2}(t)}{2}\right)I^{2}+I\left(N\beta-\gamma-\mu-\frac{\sigma^{2}(t)N^{2}}{2}\right)\right)\delta t+\sigma(t)I(N-I)\circ\delta\mathcal{B}\,.$ (2.7)
@@
$\mathfrak{S}(t,\mathcal{B},I):(\delta t,\delta\mathcal{B})\in\mathrm{T}_{(t,\mathcal{B})}\mathbb{R}^{2}\longmapsto\mathfrak{S}_{1}(t,I)\delta t+\mathfrak{S}_{2}(t,I)\delta\mathcal{B}\in\mathrm{T}_{I}\mathbb{R}\,$
@@
$\begin{gathered}\mathfrak{S}_{1}(t,I)=-\sigma^{2}(t)I^{3}+\left(-\beta+\frac{3N\sigma^{2}(t)}{2}\right)I^{2}+I\left(N\beta-\gamma-\mu-\frac{\sigma^{2}(t)N^{2}}{2}\right),\\ \mathfrak{S}_{2}(t,I)=\sigma(t)I(N-I).\end{gathered}$
@@
Recall that the theoretical utilisation of Stratonovich stochastic equations is justified by Malliavin’s Transfer Principle [45], which states that the results from the theory of ordinary differential equations remain applicable, in an analogous way, to stochastic differential equations in Stratonovich form. This principle is just a general guideline without a proof, which implies that it must be used just as a general suggestion. Nevertheless, as shown in Example 2.1, the relation to the Itô approach has to be considered too for analysing applications.
@@
for family of functions $b_{a}^{\alpha}:B\in\mathbb{R}^{\ell}\mapsto b_{a}^{\alpha}(B)\in\mathbb{R}$, with $\alpha=1,\ldots,r$ and $a=1,\ldots,\ell$, and a certain $r$-dimensional Lie algebra of vector fields on $M$ spanned by $X_{1},\ldots,X_{r}$. We call the Lie algebra $V=\langle X_{1},\ldots,X_{r}\rangle$ a Vessiot–Guldberg Lie algebra of the stochastic Lie system (2.8).
Note that the Stratonovich operator $\mathfrak{S}$ of a stochastic Lie system can be considered as an $\ell$-element family consisting of $\mathbb{R}^{\ell}$-vector fields on $M$. Each element of the family can be understood as a section of the trivial bundle ×$\mathbb{R}^{\ell}\times V\rightarrow\mathbb{R}^{\ell}$, and conversely.
@@
Let us consider a damped harmonic oscillator on $\mathrm{T}\mathbb{R}$ with adapted coordinates $\{x,y\}$ and with semi-martingale $W_{1}$ in Itô form (cf. [35, 36, 47]) given by
$\delta\begin{pmatrix}x\\ y\end{pmatrix}=\begin{pmatrix}0&1\\ -\omega_{f}^{2}&-k\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}\delta t+\begin{pmatrix}0&0\\ 0&-\sigma\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}\delta W_{1}\,.$ (2.9)
More specifically, $W_{1}$ is a one-dimensional Wiener process, which is a type of semi-martingale, while $\sigma$ is a parameter, e.g. quantifying noise. The parameters $\omega_{f}$ and $k$ are the usual parameters relative to the standard deterministic model for a dissipative harmonic oscillator
$\ddot{x}+\omega_{f}^{2}x+k\dot{x}=0\,,$
where $\omega_{f}$ is the elastic constant of the oscillator and $k$ is the friction coefficient. Let us see how this model can be considered as a linear stochastic Lie system.
Consider the vector fields on $\mathrm{T}\mathbb{R}$ given by
@@
These vector fields span a Lie algebra of vector fields $V_{D}$ isomorphic to the general Lie algebra, $\mathfrak{gl}(2,\mathbb{R})$, of $2\times 2$ matrices with real coefficients. In fact, the commutation relations in the basis (2.10) read
@@
$[X_{12},X_{21}]=X_{22}-X_{11}\,,$ $[X_{12},X_{22}]=-X_{12}\,,$ $[X_{21},X_{22}]=X_{21}\,.$
Since (2.9) is in the Itô form, expression (2.5)) shows that it can be related to a Stratonovich operator of the form
$\mathfrak{S}(t,W_{1},x,y)=\left(\begin{pmatrix}0&1\\ -\omega_{f}^{2}&-k-\sigma^{2}/2\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix},\begin{pmatrix}0&0\\ 0&-\sigma\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}\right).$
Hence, one can write
$\mathfrak{S}(t,W_{1},x,y)=\left(X_{12}-\omega_{f}^{2}X_{21}-(k+\sigma^{2}/2)X_{22},-\sigma X_{22}\right),$
which turns (2.9) into a stochastic Lie system and $V_{D}$ into an associated Vessiot–Guldberg Lie algebra. Note that more general models can be considered by assuming the coefficients $\omega_{f}$, $\sigma$, or $k$ depending on $t$ and/or $W_{1}$. One can also assume $\mathfrak{S}_{1},\mathfrak{S}_{2}$ to have as components more general linear combinations of the elements of $V_{D}$ with coefficients depending on $W_{1}$ and $t$. These models will still be stochastic Lie systems. $\triangle$
@@
It is worth noting that an Itô stochastic differential equation may seem to have the form (2.8) without being a stochastic Lie system. This shows that the stochastic case does not matches exactly the form given in classical Lie systems in the Itô realm. Let us provide an example of this from the SIS model in Example 2.1. Let us study a stochastic differential equation that is related to the deterministic SIS model for particular values $N=100$, $\beta=1/2$, $\mu=\gamma=0$, which is indeed a deterministic approximation of it. Consider the Itô stochastic differential equation
$\delta I=I(50-I/2)\delta t+I\sigma(100-I)\delta\mathcal{B}\,,$ (2.11)
for a certain $t$-dependent parameter $\sigma=\sigma(t)$, which generalises the stochastic SIS system studied in [32, p. 880]. Consider the vector fields
@@
which span a Lie algebra of vector fields with commutation constants
$[Y_{1},Y_{2}]=Y_{2},\qquad[Y_{1},Y_{3}]=2Y_{2},\qquad[Y_{2},Y_{3}]=Y_{3}.$
Moreover, the operator of the form
$\mathcal{I}(t,\mathcal{B},I)=(50Y_{2}-1/2Y_{3},100\sigma(t)Y_{2}-\sigma(t)Y_{3}),$
is related to the Vessiot–Guldberg Lie algebra $\langle Y_{1},Y_{2},Y_{3}\rangle$ and one would expected (2.11) to be a stochastic Lie system. Nevertheless, one has that (2.11) is related to the Stratonovich differential equation
$\delta I=\left(-\sigma(t)I^{3}+\left(-\frac{1}{2}+150\sigma^{2}(t)\right)I^{2}+I(50-5000\sigma^{2}(t))\right)\delta t+I\sigma(t)(100-I)\circ\delta\mathcal{B}\,,$ (2.12)
which cannot be described by the Vessiot–Guldberg Lie algebra $\langle Y_{1},Y_{2},Y_{3}\rangle$ and any other larger one (all of them are a generic point isomorphic on $\mathbb{R}$ to a Lie subalgebra of $\mathfrak{sl}_{2}$ (see [20, 30] for details). Indeed, one can prove, as shown hereafter in Section 3, that (2.12) is not a stochastic Lie system. $\triangle$
@@
We call $t$-dependent stochastic Riccati differential equation the Stratonovich stochastic differential equation on $\mathbb{R}$ with stochastic variables given by the semi-martingales $\mathcal{B}^{2},\ldots,\mathcal{B}^{\ell}:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}$ of the form
$\delta\Gamma=\left(\sum_{\alpha=0}^{2}b_{\alpha}(t)\Gamma^{\alpha}\right)\delta t+\sum_{\beta=2}^{\ell}\sum_{\alpha=0}^{2}b_{\beta\alpha}(t)\Gamma^{\alpha}\circ\delta\mathcal{B}^{\beta},\qquad\forall\Gamma\in\mathbb{R},\qquad\forall t\in\mathbb{R}.$ (2.13)
@@
$t$-Dependent stochastic Riccati differential equations are stochastic Lie systems. The following proposition is immediate.
@@
for arbitrary $t$-dependent functions $b_{\alpha}(t),b_{\mu\alpha}(t)$ with $\alpha=0,1,2$ and $\mu=2,\ldots,\ell$ is also a noise-free stochastic Riccati differential equation.
It is worth noting that the theory of Lie systems can be generalised to many different realms [20, 23]. In particular, there is the theory of foliated Lie systems [17]. This suggests us the following generalisation.
@@
where $B:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}^{\ell}$ is an $\mathbb{R}^{\ell}$-valued semi-martingale and $\mathfrak{S}(B,\Gamma):\mathrm{T}_{B}\mathbb{R}^{\ell}\to\mathrm{T}_{\Gamma}M$ is a Stratonovich operator, such that
@@
and the vector fields $\{Y_{1},\ldots,Y_{r}\}$ on $M$ span an $r$-dimensional Vessiot–Guldberg Lie algebra such that $b^{\alpha}_{j}(B,\Gamma)$, with $\alpha=1,\ldots,r$ and $j=1,\ldots,\ell$, are first integrals of the vector fields $Y_{1},\ldots,Y_{r}$ for every $B\in\mathbb{R}^{\ell}$.
It is worth noting that there are many potential applications concerning the generalisation to the stochastic realm of famous types of Lie systems such as matrix, projective Riccati equations, Bernouilli equations, and so on. As illustrated in this section, stochastic Lie systems have applications in predator-prey models, oscillator type models, et cetera [3, 35, 52]. Moreover, nonlinear stochastic differential equations are difficult to study. Notwithstanding, under certain conditions, its linearisation can describe its equilibrium properties [3]. Linear or even affine stochastic differential equations with stochastic variables related to semi-martingales are stochastic Lie systems (cf. [39]).

### v3

@@ 2 Stochastic differential equations and stochastic Lie systems
where the integrals appearing above are Stratonovich integrals. It is relevant to understand that stochastic differential equations can still be understood in the so-called Itô way, in which the general solution is also of the form (2.4), but the integrals are assumed to be Itô integrals, which is different. Which approach is used depends on the applications to be developed. It should be observed that in the interpretation of stochastic differential equations according to Itô, the symbol $\circ$ is omitted.
Unless otherwise stated, stochastic differential equations are assumed to be in the Stratonovich sense. This is motivated by the Malliavin transfer principle [50, 53], which suggests that the obtained theory will retain the standard differential theory, although it may not always be the case. In other words, Malliavin transfer principle states that “A formula which is true in the deterministic context and which has a meaning via Stratonovich stochastic calculus, is still valid, but only almost surely” [50].
Despite the Malliavin transfer principle, stochastic differential equations appear in the Itô sense in many applications [29, 56, 67]. Hence, a manner to deal with such stochastic differential equations will be studied in this work. Moreover, the analysis of such stochastic differential equations shows some differences appearing in the stochastic formalism with respect to the deterministic theory of Lie systems, which enriches the theory.
Although every Stratonovich stochastic differential equation is equivalent to another Itô stochastic differential equation, the form of both is different. Hence, one has to take care of the method employed to study a stochastic differential equation [30]. More exactly, assume that (2.2) has coefficients that do not depend on ${B}^{2},\ldots,{B}^{\ell}$, then the Itô differential equation admitting the same solutions reads (see [15, p. 137] for details)
@@
$\delta I=\left(-\sigma^{2}(t)I^{3}+\left(-\beta+\frac{3N\sigma^{2}(t)}{2}\right)I^{2}+\left(N\beta-\gamma-\mu-\frac{\sigma^{2}(t)N^{2}}{2}\right)I\right)\delta t+\sigma(t)(N-I)I\circ\delta\mathcal{B}\,.$ (2.7)
@@
$\mathfrak{S}(t,\mathcal{B},I):(\delta t,\delta\mathcal{B})\in\mathrm{T}_{(t,\mathcal{B})}\mathbb{R}^{2}\longmapsto\mathfrak{S}_{1}(t,I)\delta t+\mathfrak{S}_{2}(t,I)\circ\delta\mathcal{B}\in\mathrm{T}_{I}\mathbb{R}\,$
@@
$\begin{gathered}\mathfrak{S}_{1}(t,I)=-\sigma^{2}(t)I^{3}+\left(-\beta+\frac{3N\sigma^{2}(t)}{2}\right)I^{2}+\left(N\beta-\gamma-\mu-\frac{\sigma^{2}(t)N^{2}}{2}\right)I,\\ \mathfrak{S}_{2}(t,I)=\sigma(t)(N-I)I.\end{gathered}$
Geometrically, each component of $\mathfrak{S}$ can be understood as an $\mathbb{R}^{2}$-dependent vector field. More specifically, $\mathfrak{S}_{1}$ and $\mathfrak{S}_{2}$ can be understood as $\mathbb{R}^{2}$-dependent vector fields
$\left[-\sigma^{2}(t)I^{3}+\left(-\beta+\frac{3N\sigma^{2}(t)}{2}\right)I^{2}+\left(N\beta-\gamma-\mu-\frac{\sigma^{2}(t)N^{2}}{2}\right)I\right]\frac{\partial}{\partial\Gamma},\qquad\sigma(t)(N-I)I\frac{\partial}{\partial\Gamma}.$
@@
Equations (2.5) and the last example illustrate the central role of the Stratonovich formulation. To make the relation with the Itô formulation explicit, we state the following standard lemma [44, 58].
Lemma 2.2 (Itô–Stratonovich correction in coordinates).
Let $X=(X^{1},\dots,X^{n})$ be a Stratonovich diffusion on a smooth manifold $M$ with local coordinates $(x^{i})$, written as
$\mathrm{d}x^{i}_{t}=A^{i}(x_{t})\,\mathrm{d}t+\sum_{k=1}^{m}B^{i}_{k}(x_{t})\circ\mathrm{d}B^{k}_{t}.$
Then the corresponding Itô form is
$\mathrm{d}x^{i}_{t}=\Bigl{(}A^{i}(x_{t})+\tfrac{1}{2}\sum_{k=1}^{m}\sum_{j=1}^{n}B^{j}_{k}(x_{t})\,\frac{\partial B^{i}_{k}}{\partial x^{j}}(x_{t})\Bigr{)}\mathrm{d}t+\sum_{k=1}^{m}B^{i}_{k}(x_{t})\,\mathrm{d}B^{k}_{t}.$
Equivalently, in tensorial notation
[…]
independently of the choice of torsion-free connection $\nabla$.
The additional drift term $\tfrac{1}{2}\sum_{k=1}^{m}\nabla_{X_{k}}X_{k}$ may not belong to the finite-dimensional Lie algebra generated by $\{X_{0},\dots,X_{m}\}$. Therefore, a system that is a Lie system in the Stratonovich sense may cease to be one in the Itô formulation. This observation is important for applications, where many models are naturally written in Itô form, and motivates our systematic use of the Stratonovich framework in the theory of stochastic Lie systems.
Recall that the theoretical utilisation of Stratonovich stochastic equations is justified by Malliavin’s Transfer Principle [53], which states that the results from the theory of ordinary differential equations remain applicable, in an analogous way, to stochastic differential equations in Stratonovich form. This principle is just a general guideline without a proof, which implies that it must be used just as a general suggestion. Nevertheless, as shown in Example 2.1, the relation to the Itô approach has to be too considered for analysing applications.
Let us now turn to analysing the main type of stochastic differential equations to be studied in this work (see [48] for the pioneering work on the topic).
@@
for a family of functions $b_{a}^{\alpha}:B\in\mathbb{R}^{\ell}\mapsto b_{a}^{\alpha}(B)\in\mathbb{R}$ that are assumed to be non-anticipative, i.e. measurable with respect to the natural filtration, with $\alpha=1,\ldots,r$ and $a=1,\ldots,\ell$, and a certain $r$-dimensional Lie algebra of vector fields on $M$ spanned by $X_{1},\ldots,X_{r}$. We call the Lie algebra $V=\langle X_{1},\ldots,X_{r}\rangle$ a Vessiot–Guldberg Lie algebra of the stochastic Lie system (2.8).
As for any other Stratonovich stochastic differential equation (2.3), recall that the Stratonovich operator $\mathfrak{S}=(\mathfrak{S}_{1},\ldots,\mathfrak{S}_{\ell})$ of a stochastic Lie system can be considered as an $\ell$-element family consisting of $\mathbb{R}^{\ell}$-vector fields on $M$ of the form $\mathfrak{S}_{\alpha}:(B,\Gamma)\in\mathbb{R}^{\ell}\times M\mapsto\mathfrak{S}_{\alpha}(B,\Gamma)\in\mathrm{T}_{\Gamma}M\subset\mathrm{T}M$ for $\alpha=1,\ldots,\ell$. Moreover, each one of these $\mathbb{R}^{\ell}$-vector fields on $M$ is a linear combination with functions depending on $\mathbb{R}^{\ell}$ of a finite-dimensional Lie algebra of vector fields $X_{1},\ldots,X_{r}$. Recall that this implies that there exists constants $c_{\alpha\beta}^{\gamma}$, with $\alpha,\beta,\gamma=1,\ldots,r$, so that
$X_{\alpha},X_{\beta}]=\sum_{\gamma=1}^{r}c_{\alpha\beta}^{\gamma}X_{\gamma},\qquad\alpha,\beta=1,\ldots,r.$
It is very important to stress that $c_{\alpha,\beta}^{\gamma}$ are constants. Although a stringent condition, it is justified by the significant applications and geometric characteristics of Lie systems (refer to [23, 26] where over a two hundred related works, applications, and key authors of Lie systems are cited) and stochastic counterparts in [49] and this work. Let us say in advance that, as shown in following parts of this work, the dimension and nature of the Vessiot–Guldberg Lie algebra is related to the properties of the stochastic Lie system. In standard Lie systems, where $B$ is just the time, a solvable Vessiot–Guldberg Lie algebra ensures, for instance, that the associated Lie system can be integrated by quadratures [22].
@@
Let us consider a generalisation of a damped harmonic oscillator on $\mathrm{T}\mathbb{R}$ with a stochastic part modelled by means of a semi-martingale $W_{1}$ related to a one-dimensional Wiener process in Itô form, which retrieves as a particular cases several models in the previous literature (cf. [41, 42, 55]). In particular, consider adapted coordinates $(x,y=\dot{x})$ on $\mathrm{T}\mathbb{R}$ and the stochastic differential equation given by
$\delta\begin{pmatrix}x\\ y\end{pmatrix}=\begin{pmatrix}0&1\\ -\omega_{f}^{2}(t)&-k(t)\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}\delta t+\begin{pmatrix}0&0\\ 0&-\sigma(t)\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}\delta W_{1}\,$, (2.9)
where $\sigma(t)$ is any $t$-dependent function, e.g. quantifying noise, and the functions $\omega_{f}(t)$ and $k(t)$ are extensions to a stochastic realm of the usual functions relative to the standard deterministic model for a dissipative harmonic oscillator
$\ddot{x}+\omega_{f}^{2}(t)x+k(t)\dot{x}=0\,.$
Then, $\omega_{f}(t)$ is a $t$-dependent frequency of the oscillator (2.9) while $k(t)$ is frequently used to describe a friction-like effect.
Let us see how the model (2.9) can be considered as a linear stochastic Lie system. The first step needed to apply our formalism is to transform the previous stochastic system from an Itô into a Stratonovich one. This can be reached in a simple manner by applying the transformation equation given in (2.5). Indeed, this expression shows that (2.9) can be related to a Stratonovich operator of the form
$\mathfrak{S}(t,W_{1},x,y)=\left(\begin{pmatrix}0&1\\ -\omega_{f}^{2}(t)&-k(t)-\sigma^{2}(t)/2\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix},\begin{pmatrix}0&0\\ 0&-\sigma(t)\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}\right)$. (2.10)
It is worth recalling that many stochastic differential equations are formulated in the Itô framework.
Once the Stratonovich form has been obtained, let us show that one may apply the theory of stochastic Lie systems to this example. With this aim, one has to recall that the two components of the Stratonovich operator (2.10) are related to two $t$-dependent vector fields, corresponding to its two components, $\mathfrak{S}_{1},\mathfrak{S}_{2}$, of the form
$\mathfrak{S}_{1}^{x}\frac{\partial}{\partial x}+\mathfrak{S}_{1}^{y}\frac{\partial}{\partial y}=y\frac{\partial}{\partial x}-(\omega_{f}^{2}(t)x+[k(t)+\sigma^{2}(t)/2]y)\frac{\partial}{\partial y},\qquad\mathfrak{S}_{2}^{x}\frac{\partial}{\partial x}+\mathfrak{S}_{2}^{y}\frac{\partial}{\partial y}=-\sigma(t)y\frac{\partial}{\partial y},$
respectively. Note that both $t$-dependent vector fields can be considered as $\mathbb{R}^{2}$-dependent vector fields that have a trivial dependence on the variable $W_{1}$ in $(t,W_{1})\in\mathbb{R}^{2}$. To describe our model with the theory of stochastic Lie systems, one has to find some finite-dimensional Lie algebra of vector fields on $\mathrm{T}\mathbb{R}^{2}$, let us say $V^{D}$, such that each component of the Stratonovich operator, which can be understood as an $\mathbb{R}^{2}$-dependent vector field, becomes a linear combination with coefficients depending on $t,W_{1}$ of a basis of $V_{D}$.
In this case, the sought Lie algebra of vector fields can be obtained by considering the vector fields on $\mathrm{T}\mathbb{R}$ given by
@@
These vector fields span a four-dimensional Lie algebra of vector fields $V_{D}$ isomorphic to the general Lie algebra, $\mathfrak{gl}_{2}$, of $2\times 2$ matrices with real coefficients. In fact, the commutation relations in the basis (2.11) read
@@
$[X_{12},X_{21}]=X_{22}-X_{11}\,$, $[X_{12},X_{22}]=-X_{12}\,$, $[X_{21},X_{22}]=X_{21}\,$,
and they are equal to the commutation relations between
$M_{11}=\left[\begin{matrix}-1&0\\ 0&0\end{matrix}\right],\,\,M_{12}=\left[\begin{matrix}0&-1\\ 0&0\end{matrix}\right],\,\,M_{21}=\left[\begin{matrix}0&0\\ -1&0\end{matrix}\right],\,\,M_{22}=\left[\begin{matrix}0&0\\ 0&-1\end{matrix}\right].$
Hence, the Stratonovich operator (2.10) describing our model can be written as
$\mathfrak{S}(t,W_{1},x,y)=\left(X_{12}-\omega_{f}^{2}(t)X_{21}-(k(t)+\sigma^{2}(t)/2)X_{22},-\sigma(t)X_{22}\right).$
Thus, each component of the Stratonovich operator can be written as a linear combination with coefficients depending on $t,W_{1}$ (in particular the coefficients of our model only depend on $t$), of the vector fields of a basis of $V_{D}$. Hence, (2.9) is a stochastic Lie system and $V_{D}$ becomes an associated Vessiot–Guldberg Lie algebra.
Note that more general models can be considered as stochastic Lie systems by assuming that their Stratonovich operators take the form
$\mathfrak{S}(t,W_{1},x,y)=\left(\sum_{\alpha,\beta=1}^{2}b_{\alpha\beta}(t,W_{1})X_{\alpha\beta},\sum_{\alpha,\beta=1}^{2}b^{W}_{\alpha\beta}(t,W_{1})X_{\alpha\beta}\right),$
for arbitrary functions $b_{\alpha\beta},b^{W}_{\alpha\beta}:\mathbb{R}^{2}\rightarrow\mathbb{R}$, which also admit a Vessiot–Guldberg Lie algebra $V_{D}$. One could choose even larger Vessiot–Guldberg Lie algebras, e.g. the Lie algebra of affine vector fields on $\mathbb{R}^{2}$. Moreover, if the functions $k(t)$, $\omega_{f}(t)$ and $\sigma(t)$ take particular values, e.g. constant ones, it may happen that one could choose a smaller Vessiot–Guldberg Lie algebra. For instance, if $k(t)=\omega_{f}(t)=\sigma(t)=0$, the Stratonovich operator could be described via a Vessiot–Guldberg Lie algebra spanned by $X_{12}$. Depending on the dimension of the Vessiot–Guldberg Lie algebra, the superposition rule may depend on less or more particular solutions, making numerical methods simpler or more difficult to be applied. This fact will be exaplained in detail after the stochastic Lie theorem in Section 3.  $\triangle$
@@
It is worth noting that an Itô stochastic differential equation $\delta\Gamma=\mathfrak{S}(B,\Gamma)\delta B$ with $\mathfrak{S}(B,\Gamma)$ taking the form (2.8) may not be a stochastic Lie system. This shows that stochastic Lie systems, in the Itô framework, do not match exactly the form given in classical Lie systems. Let us provide an example of this, with practical implications, using the SIS model in Example 2.1. Let us study a stochastic differential equation $\delta I=\mathcal{I}(t,\mathcal{B},I)\delta B$ related to the deterministic SIS model for particular values $N=100$, $\beta=1/2$, $\mu=\gamma=0$, which is indeed a deterministic approximation of it. More specifically, consider the Itô stochastic differential equation
$\delta I=(50-I/2)I\delta t+\sigma(100-I)I\delta\mathcal{B}\,$, (2.12)
for a $t$-dependent parameter $\sigma=\sigma(t)$ which is not constant. This model generalises the stochastic SIS system studied in [35, p. 880]. Consider the vector fields
@@
which span a three-dimensional Lie algebra $V_{R}$ of vector fields with commutation constants
$Y_{1},Y_{2}]=Y_{2},\qquad[Y_{1},Y_{3}]=2Y_{2},\qquad[Y_{2},Y_{3}]=Y_{3}.$
In fact, this Lie algebra is isomorphic to $\mathfrak{sl}_{2}$, namely the matrix Lie algebra of traceless $2\times 2$ matrices with real coefficients. Moreover, the operator $\mathcal{I}$ is such that its components can be written as linear combinations with $t$-dependent coefficients of vector fields of $V_{R}$ in the form
$\mathcal{I}(t,\mathcal{B},I)=(50Y_{2}-1/2Y_{3},100\sigma(t)Y_{2}-\sigma(t)Y_{3}).$
Hence, one notes that (2.12) looks like a stochastic Lie system, but we have to recall that the form (2.8) must appear in the Stratonovich form of our stochastic differential equations. Nevertheless, one has that (2.12) is a stochastic differential equation in the Itô form related to the Stratonovich stochastic differential equation $\delta I=\mathfrak{S}(t,\mathcal{B},I)\circ\delta B$ of the form
$\delta I=\left(-\sigma(t)I^{3}+\left(-\frac{1}{2}+150\sigma^{2}(t)\right)I^{2}+(50-5000\sigma^{2}(t))I\right)\delta t+\sigma(t)(100-I)I\circ\delta\mathcal{B}\,.$ (2.13)
The problem is that for different values of $\sigma(t)$, which is not constant by assumption, the first component of $\mathfrak{S}(t,\mathcal{B},I)$, namely
$\sigma(t)I^{3}+\left(-\frac{1}{2}+150\sigma^{2}(t)\right)I^{2}+(50-5000\sigma^{2}(t))I$
is a $t$-dependent vector field whose values at different values of $t\in\mathbb{R}$ generate the vector space $E=\langle Z_{1}=I\partial/\partial I,Z_{2}=I^{2}\partial/\partial I,Z_{3}=I^{3}\partial/\partial I\rangle$, which cannot be described by the Vessiot–Guldberg Lie algebra $\langle Y_{1},Y_{2},Y_{3}\rangle$ nor by any other. Indeed, the elements of $E$ cannot be written as a linear combination of elements of a finite-dimensional Lie algebra. In fact, the successive Lie brackets
$Z_{2},Z_{3}]=I^{4}\frac{\partial}{\partial I},\qquad[Z_{2},[Z_{2},Z_{3}]]=2I^{5}\frac{\partial}{\partial I},\qquad[Z_{2},[Z_{2},[Z_{2},Z_{3}]]]=3\cdot 2I^{6}\frac{\partial}{\partial I},\qquad\ldots$
span an infinite family of linearly independent vector fields on $\mathbb{R}$ that must be included in any Lie algebra containing $E$.  $\triangle$
@@
We call a $t$-dependent stochastic Riccati differential equation the Stratonovich stochastic differential equation on $\mathbb{R}$ with stochastic variables given by the semi-martingales $\mathcal{B}^{2},\ldots,\mathcal{B}^{\ell}:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}$ of the form
$\delta\Gamma=\left(\sum_{\alpha=0}^{2}b_{\alpha}(t)\Gamma^{\alpha}\right)\delta t+\sum_{\beta=2}^{\ell}\sum_{\alpha=0}^{2}b_{\beta\alpha}(t)\Gamma^{\alpha}\circ\delta\mathcal{B}^{\beta},\qquad\forall\Gamma\in\mathbb{R},\qquad\forall t\in\mathbb{R},$ (2.14)
@@
Note that $t$-dependent stochastic Riccati differential equations are stochastic Lie systems. The following proposition is immediate.
@@
for arbitrary $t$-dependent functions $b_{\alpha}(t),b_{\mu\alpha}(t)$ with $\alpha=0,1,2$ and $\mu=2,\ldots,\ell$ is also a $t$-dependent stochastic Riccati differential equation.
It is worth noting that the theory of Lie systems can be generalised to many different realms [23, 26]. In particular, there is the theory of foliated Lie systems [18]. This suggests the following generalisation.
@@
where $B:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}^{\ell}$ is an $\mathbb{R}^{\ell}$-valued semi-martingale and $\mathfrak{S}(B,\Gamma):\mathrm{T}_{B}\mathbb{R}^{\ell}\to\mathrm{T}_{\Gamma}M$ is a Stratonovich operator such that
@@
and the vector fields $\{Y_{1},\ldots,Y_{r}\}$ on $M$ span an $r$-dimensional Vessiot–Guldberg Lie algebra such that $b^{\alpha}_{j}(B,\Gamma)$, with $\alpha=1,\ldots,r$ and $j=1,\ldots,\ell$, are first integrals of the vector fields $Y_{1},\ldots,Y_{r}$ for every fixed $B\in\mathbb{R}^{\ell}$.
There are many potential applications concerning the generalisation to the stochastic realm of famous types of Lie systems such as matrix, projective Riccati equations, Bernouilli equations, and so on. As illustrated in this work, stochastic Lie systems have applications in predator-prey models, oscillator type models, et cetera [5, 41, 64]. Moreover, nonlinear stochastic differential equations are difficult to study. Notwithstanding, under certain conditions, their linearisation can describe their equilibrium properties [5]. Linear or even affine stochastic differential equations with stochastic variables related to semi-martingales are stochastic Lie systems (cf. [48]).